OpenStudy (anonymous):

(9,15),(-12,3) write the two euqation of a line given two points

OpenStudy (anonymous):

equation*

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

First Find the Slope \[m= \frac{ (y2-y1) }{ (x2-x1) }\] the use Two point form: \[y-y1 = \frac{ (y2-y1) }{ (x2-x1) } ( x-x1)\]

OpenStudy (anonymous):

yes i noe how to fin the slope but not the equation its confusing

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

the slope is already -12/-21

jimthompson5910 (jim_thompson5910):

keep going and reduce that

OpenStudy (anonymous):

can u doo it for me so i could understand ?

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (3-15)/(-12-9) m = (-12)/(-21) m = 4/7

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

I'm finding the slope

jimthompson5910 (jim_thompson5910):

(-12)/(-21) reduces to 4/7

jimthompson5910 (jim_thompson5910):

so the slope is 4/7

OpenStudy (anonymous):

yes i noe but i already noe how to find the slope i need help making the equation y=mx+b

jimthompson5910 (jim_thompson5910):

y = mx+b y = 4/7x + b ... Plug in m = 4/7 15 = (4/7)*(9) + b ... Plug in x = 9 and y = 15

jimthompson5910 (jim_thompson5910):

now solve for b